solve each of the following systmes by substitution.
4x-12y=5
-x+3y=-1

Answer by venugopalramana(3286)

solve each of the following systmes by substitution.
4x-12y=5................I
-x+3y=-1...............II.....X=3Y+1..PUT THIS IN EQN.I
4(3Y+1)-12Y=5
12Y+4-12Y=5
4=5...NOT POSSIBLE
HENCE EQNS. ARE INCONSISTENT AND THERE IS NO SOLUTION